Express the following surfaces in spherical coordinates:


y² +z² -x² = 1


1
Expert's answer
2022-01-24T02:27:31-0500

In shperical coordinates:


"x = r\cos \phi \sin \theta\\
y = r\sin \phi \sin\theta\\
z = r\cos \theta"

Substituting into the equation, obtain:

"r^2\sin^2 \phi \sin^2\theta +r^2\cos^2\theta-r^2\cos^2 \phi \sin^2\theta=1\\
r^2\sin^2\theta(\sin^2 \phi -\cos^2 \phi )+r^2\cos^2\theta=1\\
r^2\cos^2\theta - r^2\sin^2\theta\cos2\phi=1"

Answer. "r^2\cos^2\theta - r^2\sin^2\theta\cos2\phi=1".